# UniGAP: A Universal and Adaptive Graph Upsampling Approach to Mitigate Over-Smoothing in Node Classification Tasks

### v1

@@ 3.3.2. Multi-View Condensation
(7) $\hat{T}=\mathcal{F}_{\omega}(T)=\operatorname{TMM}({T})=\sigma(w_{\text{traj}}\odot T)\oplus T)W_{\text{channel}}$
@@
where $\operatorname{Tr}(\cdot)$ denotes a shallow Transformer encoder, and $\mathcal{R}$ represents an attention-based readout function.
@@
3.3.3. Adaptive Graph Upsampler
After obtaining condensed trajectory features $\hat{T}$, we use them to compute the upsampling probability for each existing edge to decelerate message passing, thereby mitigating the over-smoothing problem (theoretical justification can be found in Section 3.5.1). Subsequently, we employ a differentiable sampling method to generate an augmented graph. Specifically, we compute the upsampling logits as follows:
(9) $P_{ij}=\mathcal{U}_{\phi}(\hat{T})={W}_{\text{up}}\left[\hat{{T}}_{i}\|\hat{{T}}_{j}\right]\in\mathbb{R}^{|\mathcal{E}|\times 2},\forall(i,j)\in\mathcal{E}$
where $P_{ij}$ represents the upsampling logits for the edge between nodes $i$ and $j$, ${W}_{\text{up}}$ is a learnable weight matrix, and $\left[\hat{{T}}_{i}\|\hat{{T}}_{j}\right]$ denotes the concatenation of the condensed trajectory features of nodes $i$ and $j$.
Then we employ Gumbel-Softmax (Jang et al., 2017) with additional tricks to derive the upsampling mask and generate the augmented graph $\hat{G}$:
@@ 3.3.3. Adaptive Graph Upsampler
(10) $M=\mathds{1}_{\hat{P}[:,0]\leq\hat{P}[:,1]}-\tilde{{P}}+\hat{P}\in\mathbb{R}^{|\mathcal{E}|}$
where $\epsilon\in\mathbb{R}^{|\mathcal{E}|\times 2}$ represents Gumbel noise used to introduce diversity into the sampling process (exploration), preventing the selection of edges solely based on highest probabilities and ensuring differentiability. $\hat{P}$ denotes the upsampling probability, and $\mathds{1}_{\hat{P}[:,0]\leq\hat{P}[:,1]}$ is an indicator function that returns 1 if $\hat{P}[:,0]\leq\hat{P}[:,1]$. $\tilde{P}$ results from truncating the gradient of $\hat{P}$. This approach achieves two objectives: (1) ensures the output is exactly one-hot (by adding and then subtracting a soft value), and (2) maintains the gradient equal to the soft gradient (by keeping the original gradients). Lastly, $M$ represents an active mask used to determine which edges will have a new node inserted. Specifically, if the mask for edge $i\to j$ is 1, a new node $k$ is inserted between nodes $i$ and $j$. The original edge $i\to j$ is then masked and replaced by new edges $i\to k$ and $k\to j$. The initial feature vector of node $k$ can be set to a zero vector, the average of nodes $i$ and $j$, or an adaptive sum:
@@
Then, we obtain the augmented graph $\hat{{G}}=(\hat{\mathcal{V}},\hat{\mathcal{E}},\hat{X})$, which will be used for downstream tasks to mitigate over-smoothing and improve overall performance.
3.3.4. Model Training
To optimize the augmented graph, we will utilize it in downstream tasks such as node classification. Let the downstream model be denoted as $\mathcal{D}_{\psi}(\cdot)$. The augmented graph will be fed into this model to compute the downstream loss $\mathcal{L}_{\text{down}}$. Additionally, we incorporate over-smoothing metrics, such as Mean Average Distance (MAD) (Chen et al., 2020b) and Dirichlet Energy (Rusch et al., 2023), as regularization terms $\mathcal{L}_{\text{smooth}}$ to ensure that UniGAP explicitly focuses on mitigating the over-smoothing issue. The total loss is defined as:
@@ 3.3.4. Model Training
$=-\mathbb{E}_{v\in\hat{\mathcal{V}}}y\log p(\tilde{y}\mid v,\mathcal{N}_{v})-\beta\operatorname{MAD}(\hat{G})$
where $\tilde{y}$ represents the prediction of the downstream model, $\mathcal{N}_{v}$ denotes the set of neighboring nodes associated with node $v$ used in the downstream model, and $\beta$ is the regularization coefficient that controls the impact of the smoothing metric. We empirically set $\beta$ to 1 because a larger value would negatively impact downstream performance. The formula of MAD is represnted as:
(13) $\operatorname{MAD}(X)=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}_{i}}1-\frac{X_{i}^{T}X_{j}}{||X_{i}||||X_{j}||}$
Through optimizing the downstream loss, all parameters in our framework will be updated in an end-to-end manner through backpropagation. After warming up the downstream model, it will be used to compute trajectories instead of relying solely on the trajectory precomputation module mentioned in Section 3.3.1. The model training process continues until convergence, aiming to obtain the optimal upsampled graph and model parameters.
@@ 3.4. UniGAP as A Universal Framework
UniGAP is a universal plug-in that can be integrated with various GNN models, as demonstrated in Section 4.2.1. Additionally, by combining different components and strategies, UniGAP can encompass most existing graph upsampling techniques (Chen et al., 2020a; Azabou et al., 2023; Wu et al., 2021). In this subsection, we will discuss how these techniques can be considered as variants of our approach.
HalfHop (Azabou et al., 2023). Consider the Trajectory Computation initiated with a stochastic vector, with Multi-View Condensation and Graph Upsampler frozen with random parameters. The upsampling of nodes then becomes a stochastic process with probability $p$ based on these parameters, making UniGAP function as a randomized heuristic upsampler, equivalent to HalfHop.
AdaEdge (Chen et al., 2020a). Let the Trajectory Computation contain only node embeddings from the last layer of the downstream GNN, with the Multi-View Condensation omitted. The module computing the upsampling probability for each edge in the Adaptive Graph Upsampler can be considered a binary classifier, with the output indicating whether the edge is intra-class or inter-class. This allows UniGAP to determine edge types based on the last layer’s output, adaptively increasing intra-class connections and decreasing inter-class connections, similar to AdaEdge.
GraphMixup (Wu et al., 2021). Using Non-parametric Message Passing as the Trajectory Precomputation Module and a self-supervised MVC Encoder to learn disentangled semantic features, and constraining the Graph Upsampler to generate only synthetic minority nodes with a GNN classifier updating the upsampling scale via a reinforcement learning agent, makes the UniGAP framework equivalent to GraphMixup, improving class-imbalanced node classification tasks.
3.5. Discussion
3.5.1. Why Does UniGAP Work?
Equipped with UniGAP, the optimized upsampled graph can not only enhance downstream performance but also mitigate the over-smoothing problem. In this subsection, we will provide theoretical explanations by answering the following questions:
RQ1: How does UniGAP learn in concert with downstream tasks to get the optimal upsampling graph structure?
RQ2: How does UniGAP mitigate the over-smoothing problem?
Closed-Form Solution (RQ1). To simplify the analysis, we assume the GNN model as linear GNN with mean aggregation (Keriven, 2022). Therefore, the $k$-th hop node features at $e$-th epoch are:
(14) $H^{(k)}_{e}$ $=\hat{A}^{k}_{e}X=\operatorname{UniGAP}(H^{(k)}_{e-1},\cdots,H^{(0)}_{e-1},\theta_{u})^{k}X:=f^{k}(A,\theta_{u},X)$
@@ 3.5.1. Why Does UniGAP Work?
(15) $\langle\hat{\theta},\hat{\theta}_{u}\rangle_{e}$ $=\operatorname*{arg\,min}_{\langle\theta,\theta_{u}\rangle}\frac{1}{2n}||Y-H^{(k)}_{e}\theta||^{2}+\lambda||\theta||^{2}+\gamma||\theta_{u}||^{2}$
$\quad\quad-\beta\operatorname{MAD}(H^{(k)}_{e})$
@@
$\quad\quad\gamma^{-1}\left(\frac{1}{2n}\left(\frac{\partial f^{k}}{\partial\theta_{u}}\right)^{T}(Y-f^{k}\hat{\theta})\hat{\theta}-\frac{\beta}{2}\frac{\partial}{\partial\theta_{u}}\text{MAD}(f^{k})\right)\Bigg{\rangle}$
As shown in Equation 14,15, UniGAP can learn the message passing dynamics and adaptivly change the nodes’ receptive filed, finally reach a optimal structural representation for downstream task. The derivation and more illustrations can be found in Appendix B.
Slower Smoothing (RQ2). We find that UniGAP can decelerate the message passing process by slowing down the smooth rate of the covariance of node features during message passing. The following Lemma and Theorem shows the approximated covariance of the node features after $k$ rounds of message passing with or without UniGAP.
@@
(16) $\mathcal{R}^{(k)}_{\text{UniGAP}}\simeq\mathcal{R}_{\text{reg.}}(\frac{1}{2}A^{k-1}(I+((1-p)I+pA)^{2})\Sigma)$
Here $\Sigma^{(k)}_{\text{UniGAP}}$ is the approximated covariance of the node features, $p=\mathbb{E}(\hat{P}[:,0])$ and $\hat{P}$ denotes the upsampling probability at the optimal graph structure. The rate of smoothing of MP with UniGAP is $k-1$ due to the $A^{k-1}\Sigma$ operator. The complete proof can be found in Appendix B.
With Lemma 3.2 and 3.3, we can naturally get the Theorem:
@@
By slowing down the smoothing rate of the covariance of node features during message passing, UniGAP reduces the smoothing rate from $2k$ to $k-1$.
3.5.2. Complexity Analysis
@@
Time Complexity. Our framework comprises four primary components: trajectory computation, multi-view condensation, adaptive graph upsampler, and downstream model. The trajectory precomputation module involves complexities of $\mathcal{O}(1)$, $\mathcal{O}(LEd)$, or $\mathcal{O}(LEd+LNd^{2})$ depending on the initialization method: zero initialization, non-parametric message passing, or pre-trained GNN. The MVC Condensation module (Trajectory-MLP-Mixer) operates with a time complexity of $\mathcal{O}(LN+Nd^{2})$, while the adaptive graph upsampler computes with a complexity of $O(Ed)$ to determine upsampling probabilities for adjacent node pairs. The time complexity of downstream model, e.g., GNN, is $\mathcal{O}(LEd+LNd^{2})$. Consequently, the total time complexity of our method is $O(LN+LEd+LNd^{2})$, comparable to traditional L-layer GNNs, i.e., $\mathcal{O}(LEd+LNd^{2})$.
@@ 4.1. Experimental Setup
4.1.1. Datasets
We evaluated our method on 8 node classification benchmark datasets, as shown in Table 1, comprising 4 homophilic and 4 heterophilic datasets. More detailed information is provided in Appendix A.
4.1.2. Evaluation Protocols
To compare UniGAP with baselines, we use public split setting for all datasets, and report the mean accuracy along with the standard deviation across twenty different random seeds. UniGAP is integrated as a plug-in method with various SOTA GNN models to evaluate its effectiveness. Besides, we report the MAD metric to observe the phenomenon of over-smoothing as the number of layers increases.
4.1.3. Baselines
The baselines primarily consist of four categories: (1) traditional GNN methods such as GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), and GraphSAGE (Hamilton et al., 2017), (2) advanced GNN methods including GCNII (Chen et al., 2020c), GGCN (Yan et al., 2021), H2GCN (Zhu et al., 2020b), GPRGNN (Chien et al., 2021), and AERO-GNN (Lee et al., 2023), (3) Graph Transformer methods like GT (Dwivedi and Bresson, 2020), GraphGPS (Rampášek et al., 2023), and NAGphormer (Chen et al., 2023), and (4) other graph upsampling methods, i.e., HalfHop. We integrate UniGAP and HalfHop with various GNN models for evaluation.
4.2. Experimental Results
4.2.1. Overall Performance Comparison (RQ1)
For heterophilic datasets, UniGAP significantly enhances traditional GNN models, which heavily rely on the “homophily” assumption. For example, UniGAP achieves an average absolute improvement on four heterophilic datasets of 10.0% on GCN, 13.3% on GAT, and 5.0% on GraphSAGE, respectively. It also enhances other advanced GNN models, with absolute improvements such as 2.3% on GCNII, 1.1% on H2GCN, and 2.9% on GPRGNN. Compared to HalfHop, UniGAP surpasses it in all heterophilic datasets with over 3.9% absolute improvement with GCN, demonstrating the effectiveness of adaptive graph upsampling in identifying the optimal graph structure to enhance the downstream model’s performance.

### v2

@@ 3.2. Multi-View Condensation
(6) $\hat{T}=\mathcal{F}_{\omega}(T)=\operatorname{TMM}({T})=\sigma(w_{\text{traj}}\odot T)\oplus TW_{\text{channel}}$
@@
where $\operatorname{Tr}(\cdot)$ denotes a shallow Transformer encoder and $\mathcal{R}$ represents an attention-based readout function.
@@
3.3. Adaptive Graph Upsampler
After obtaining condensed trajectory features $\hat{T}$, we use them to compute the upsampling probability for each existing edge to decelerate message passing, thus mitigating the over-smoothing problem (theoretical justification can be found in Section 4.2). Subsequently, we employ a differentiable sampling method to generate an augmented graph. Specifically, we compute the upsampling logits as follows:
(8) $P_{ij}={W}_{\text{up}}\left[\hat{{T}}_{i}\|\hat{{T}}_{j}\right],\,\forall(i,j)\in\mathcal{E},$
where $P_{ij}\in\mathbb{R}^{2}$ represents the upsampling logits for the edge between nodes $i$ and $j$, ${W}_{\text{up}}$ is a learnable weight matrix, and $\left[\hat{{T}}_{i}\|\hat{{T}}_{j}\right]$ denotes the concatenation of the condensed trajectory features of nodes $i$ and $j$.
Then we employ straight-through Gumbel-Softmax (Jang et al., 2017; Bengio et al., 2013) to derive the upsampling mask and generate the augmented graph $\hat{G}$:
@@ 3.3. Adaptive Graph Upsampler
(9) $M=\mathds{1}_{\hat{P}[:,0]\leq\hat{P}[:,1]}-\operatorname{sg}(\hat{P})+\hat{P}\in\mathbb{R}^{|\mathcal{E}|}$
where $\epsilon\in\mathbb{R}^{|\mathcal{E}|\times 2}$ represents Gumbel noise used to introduce diversity into the sampling process (exploration), preventing the selection of edges solely based on the highest probabilities and ensuring the differentiability. $\hat{P}$ denotes the upsampling probability, and $\mathds{1}_{\hat{P}[:,0]\leq\hat{P}[:,1]}$ is an indicator function that returns 1 if $\hat{P}[:,0]\leq\hat{P}[:,1]$. $\operatorname{sg(\cdot)}$ denotes the stop gradient operator. This approach achieves two objectives: (1) ensures that the output is exactly one-hot (by adding and then subtracting a soft value), and (2) maintains the gradient equal to the soft gradient (by keeping the original gradients). $M$ represents an active mask used to determine which edges will have a new node inserted. Specifically, if the mask for edge $i\to j$ is 1, a new node $k$ is inserted between nodes $i$ and $j$. Otherwise, the original edge $i\to j$ is then masked and replaced by new edges $i\to k$ and $k\to j$. The initial feature vector of node $k$ can be set to an adaptive sum:
@@
[…]
3.4. Model Training
To jointly optimize the augmented graph and the downstream model $\mathcal{D}_{\psi}(\cdot)$ through downstream tasks, such as node classification, the augmented graph will be input into the downstream model to compute the downstream loss $\mathcal{L}_{\text{down}}$. Additionally, we incorporate over-smoothing metrics, such as Mean Average Distance (MAD) (Chen et al., 2020b) and Dirichlet Energy (Rusch et al., 2023), as regularization terms $\mathcal{L}_{\text{smooth}}$ to ensure that UniGAP explicitly focuses on mitigating the over-smoothing issue. The total loss is defined as:
@@ 3.4. Model Training
[…]
where $\tilde{y}$ represents the prediction of the downstream model, $\mathcal{N}_{v}$ denotes the set of neighboring nodes associated with node $v$ used in the downstream model and $\beta$ is the regularization coefficient that controls the impact of the smoothing metric. More analysis about $\beta$ is detailed in the Appendix A.3. In the experiments, we used the same settings for all baselines to ensure that the comparisons are fair. We empirically set $\beta$ to 1 because a larger value would negatively impact downstream performance. The formula of MAD is represented as:
[…]
Through optimizing the downstream loss, all parameters in our framework will be updated in an end-to-end manner through backpropagation. After warming up the downstream model, it will be used to compute trajectories instead of relying solely on the trajectory precomputation module mentioned in Section 3.1. The model training process continues until convergence, aiming to obtain the optimal upsampled graph and model parameters.
4. Theoretical Analysis and Discussion
In this section, we will first explain UniGAP as a universal framework in Section 4.1. Then, we will provide a theoretical analysis of why UniGAP works in Section 4.2. Last, we will present the complexity analysis of UniGAP in Section 4.3.
@@ 4.1. UniGAP as A Universal Framework
UniGAP is a universal plug-in that can be integrated with various GNN models, as demonstrated in Section 5.2. Additionally, UniGAP is a fully differentiable trajectory-based framework whose components, including the Trajectory Precomputation module, the MVC encoder, and the graph modification strategy, are all modular and can be replaced in different ways. In this subsection, we discuss its potential conceptual relationships with methods such as HalfHop and AdaEdge (Azabou et al., 2023; Chen et al., 2020a).
HalfHop (Azabou et al., 2023). With Multi-View Condensation and Graph Upsampler frozen with random parameters. The upsampling of nodes then becomes a stochastic process with probability $p$ based on these parameters, making UniGAP function as a randomized heuristic upsampler, similar to HalfHop.
AdaEdge (Chen et al., 2020a). With Trajectory Computation contain only the last layer’s embedding of the downstream GNN, the upsampling probability for each edge in the Adaptive Graph Upsampler can be considered a binary classifier of intra-class or inter-class. This allows UniGAP adaptively increasing intra-class connections and decreasing inter-class connections, similar to AdaEdge.
4.2. Why Does UniGAP Work?
Equipped with UniGAP, the optimized upsampled graph can mitigate the over-smoothing problem, thus enhancing downstream performance. In this subsection, we will provide theoretical explanations of: (1) UniGAP can learn in concert with downstream tasks to get the optimal upsampling graph structure. (2) UniGAP is able to slow the over-smoothing problem. These analyses are based on a simplified setting based on linear GNNs with mean aggregation, following prior theoretical work on over-smoothing. The goal of these analyses is to provide intuition about the smoothing dynamics induced by graph upsampling, rather than to exactly characterize the behavior of the non-linear architectures used in our experiments.
Closed-Form Solution. To simplify the analysis, we assume the GNN model as linear GNN with mean aggregation (Keriven, 2022). Therefore, the $k$-th hop node features at $e$-th epoch are:
(13) $H^{(k)}_{e}$ $=\hat{A}^{k}_{e}X$
$=\operatorname{UniGAP}(H^{(k)}_{e-1},\cdots,H^{(0)}_{e-1},\theta_{u})^{k}X$
$=f^{k}(A,\theta_{u},X)$
@@ 4.2. Why Does UniGAP Work?
[…]
$\quad-\beta\operatorname{MAD}(H^{(k)}_{e})$
@@
$\quad\gamma^{-1}\left(\frac{1}{2n}\left(\frac{\partial f^{k}}{\partial\theta_{u}}\right)^{T}(Y-f^{k}\hat{\theta})\hat{\theta}-\frac{\beta}{2}\frac{\partial}{\partial\theta_{u}}\text{MAD}(f^{k})\right)\Bigg\rangle$
where the subscript $e$ means the $e$-th epoch. As shown in Equation 13, 14, UniGAP can learn the message passing dynamics and adaptivly change the nodes’ receptive filed, finally reach a optimal structural representation for downstream task. The derivation and more illustrations can be found in Appendix B.
Slower Smoothing. We find that UniGAP can decelerate the message passing process by slowing down the smooth rate of the covariance of node features during message passing. The following Lemma and Theorem shows the approximated covariance of the node features after $k$ rounds of message passing with or without UniGAP.
Definition 0.
Following  (Azabou et al., 2023; Keriven, 2022), for a symmetric positive semi-definite matrix $S\in\mathbb{R}^{d\times d}$, we define the risk function
(15) $\mathcal{R}_{\text{reg.}}(S)\overset{def_{.}}{=}(\Sigma^{\frac{1}{2}}\beta^{*})^{T}K(\Sigma^{\frac{1}{2}}\beta^{*})\in\mathbb{R}_{+}$
where $K=(Id-S^{\frac{1}{2}}M(\gamma Id+M^{T}SM)^{-1}M^{T}S^{\frac{1}{2}})^{2}$, $\Sigma$ is the latent model covariance, $M$ is the projection matrix, $\beta^{*}$ are the true model parameters, and $\gamma$ is the ridge penalty in the least-squares estimator.
@@
(16) $\mathcal{R}_{\text{reg.}}(\Sigma^{(k)}_{\text{UniGAP}})\simeq\mathcal{R}_{\text{reg.}}(\frac{1}{2}A^{k-1}(I+((1-p)I+pA)^{2})\Sigma)$
Here $\Sigma^{(k)}_{\text{UniGAP}}$ is the approximated covariance of the node features, $p=\mathbb{E}(\hat{P}[:,0])$ and $\hat{P}$ denotes the upsampling probability at the optimal graph structure. The rate of smoothing of MP with UniGAP is k1k-1 due to the $A^{k-1}\Sigma$ operator. The complete proof can be found in Appendix B.
With Lemma 4.3 and 4.4, we can naturally get the Theorem:
@@
By slowing down the smoothing rate of the covariance of node features during message passing from $2k$ to k1k-1, UniGAP roughly half the over-smoothing problem.
4.3. Complexity Analysis
@@
Time Complexity. Our framework comprises four primary components: trajectory computation, multi-view condensation, adaptive graph upsampler, and downstream model. The trajectory precomputation module involves complexities of $\mathcal{O}(1)$, $\mathcal{O}(LEd)$, or $\mathcal{O}(LEd+LNd^{2})$ depending on the initialization method: zero initialization, non-parametric message passing, or pre-trained GNN. The MVC Condensation module (Trajectory-MLP-Mixer) operates with a time complexity of $\mathcal{O}(LN+Nd^{2})$, while the adaptive graph upsampler computes with a complexity of $O(Ed)$ to determine upsampling probabilities for adjacent node pairs. The time complexity of downstream model, e.g., GNN, is $\mathcal{O}(LEd+LNd^{2})$. Consequently, the total time complexity of our method is $\mathcal{O}(LN+LEd+LNd^{2})$, comparable to traditional L-layer GNNs, i.e., $\mathcal{O}(LEd+LNd^{2})$.
@@ 5.1. Experimental Setup
Datasets. We evaluated our method on 11 node classification benchmark datasets, as shown in Table 1 and Table 2, comprising 4 homophilic (1 large-scale) and 7 heterophilic datasets (4 large-scale). More detailed information is provided in Appendix A.
Evaluation Protocols. To compare UniGAP with baselines, we use public split setting for all datasets, and report the mean accuracy along with the standard deviation across twenty different random seeds. UniGAP is integrated as a plug-in method with various SOTA GNN models to evaluate its effectiveness. Besides, we report the MAD metric to observe the phenomenon of over-smoothing as the number of layers increases.
Baselines. The baselines primarily consist of four categories: (1) traditional GNN methods such as GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), and GraphSAGE (Hamilton et al., 2017); (2) advanced GNN methods including GCNII (Chen et al., 2020c), GGCN (Yan et al., 2021), H2GCN (Zhu et al., 2020a), GPRGNN (Chien et al., 2021), AERO-GNN (Lee et al., 2023), DRGCN (Attali et al., 2024) and MTGCN (Pei et al., 2024); (3) Graph Transformer methods like GT (Dwivedi and Bresson, 2020), GraphGPS (rampášek2023recipe), and NAGphormer (Chen et al., 2023); and (4) other graph upsampling methods, i.e., AdaEdge, HalfHop and non-upsampling methods such as DropEdge (Rong et al., 2020). We integrate UniGAP, DropEdge, HalfHop and AdaEdge with various GNN models for evaluation.
5.2. Overall Performance Comparison (RQ1)
5.2.1. Experiments on Small and Medium-scale Datasets
For heterophilic datasets, UniGAP significantly enhances traditional GNN models, which heavily rely on the “homophily” assumption. For example, UniGAP achieves an average absolute improvement on four heterophilic datasets of 10.0% on GCN, 13.3% on GAT, and 5.0% on GraphSAGE, respectively. It also enhances other advanced GNN models, with absolute improvements such as 2.3% on GCNII, 1.1% on H2GCN, and 2.9% on GPRGNN. Compared to HalfHop, UniGAP surpasses it in all heterophilic datasets with over 3.9% average absolute improvement with GCN, demonstrating the effectiveness of adaptive graph upsampling in identifying the optimal graph structure to enhance the downstream model’s performance.